\begin{thebibliography}{10}

\bibitem{ber74}
D.~P. Bertsekas.
\newblock Necessary and sufficient conditions for existence of an optimal
  portfolio.
\newblock {\em Journal of Economic Theory}, 8(2):235--247, 1974.

\bibitem{bs78}
D.~P. Bertsekas and S.~E. Shreve.
\newblock {\em Stochastic optimal control}, volume 139 of {\em Mathematics in
  Science and Engineering}.
\newblock Academic Press Inc. [Harcourt Brace Jovanovich Publishers], New York,
  1978.
\newblock The discrete time case.

\bibitem{ber76}
Dimitri~P. Bertsekas.
\newblock {\em Dynamic programming and stochastic control}.
\newblock Academic Press [Harcourt Brace Jovanovich Publishers], New York,
  1976.
\newblock Mathematics in Science and Engineering, 125.

\bibitem{bis73}
Jean-Michel Bismut.
\newblock Int\'egrales convexes et probabilit\'es.
\newblock {\em J. Math. Anal. Appl.}, 42:639--673, 1973.

\bibitem{cv77}
C.~Castaing and M.~Valadier.
\newblock {\em Convex analysis and measurable multifunctions}.
\newblock Springer-Verlag, Berlin, 1977.
\newblock Lecture Notes in Mathematics, Vol. 580.

\bibitem{chs3}
C.~Choirat, C.~Hess, and R.~Seri.
\newblock A functional version of the birkhoff ergodic theorem for a normal
  integrand: A variational approach.
\newblock {\em Annals of probability}, 31(1):63--92, 2003.

\bibitem{ds6}
F.~Delbaen and W.~Schachermayer.
\newblock {\em The Mathematics of Arbitrage}.
\newblock Springer Finance. Springer-Verlag, Berlin Heidelberg, 2006.

\bibitem{de76}
E.~B. Dynkin and I.~V. Evstigneev.
\newblock Regular conditional expectation of correspondences.
\newblock {\em Teor. Verojatnost. i Primenen.}, 21(2):334--347, 1976.

\bibitem{evs76}
I.~V. Evstigneev.
\newblock Measurable selection and dynamic programming.
\newblock {\em Math. Oper. Res.}, 1(3):267--272, 1976.

\bibitem{fs4}
H.~F{\"o}llmer and A.~Schied.
\newblock {\em Stochastic finance}, volume~27 of {\em de Gruyter Studies in
  Mathematics}.
\newblock Walter de Gruyter \& Co., Berlin, extended edition, 2004.
\newblock An introduction in discrete time.

\bibitem{kab99}
Y.~M. Kabanov.
\newblock Hedging and liquidation under transaction costs in currency markets.
\newblock {\em Finance and Stochastics}, 3(2):237--248, 1999.

\bibitem{ks9}
Y.~M. Kabanov and M.~Safarian.
\newblock {\em Markets with transaction costs}.
\newblock Springer Finance. Springer-Verlag, Berlin, 2009.
\newblock Mathematical theory.

\bibitem{kz3}
I.~Karatzas and {\u{Z}itkovi\'c}~G.
\newblock Optimal consumption from investment and random endowment in
  incomplete semimartingale markets.
\newblock {\em The Annals of Probability}, 31(4):1821–1858, 2003.

\bibitem{kn76}
J.~L. Kelley and I.~Namioka.
\newblock {\em Linear topological spaces}.
\newblock Springer-Verlag, New York, 1976.
\newblock With the collaboration of W. F. Donoghue, Jr., Kenneth R. Lucas, B.
  J. Pettis, Ebbe Thue Poulsen, G. Baley Price, Wendy Robertson, W. R. Scott,
  and Kennan T. Smith, Second corrected printing, Graduate Texts in
  Mathematics, No. 36.

\bibitem{ks99}
D.~Kramkov and W.~Schachermayer.
\newblock The condition on the asymptotic elasticity of utility functions and
  optimal investment in incomplete markets.
\newblock {\em Annals of Applied Probability}, 9(3):904--950, 1999.

\bibitem{pen11}
T.~Pennanen.
\newblock Arbitrage and deflators in illiquid markets.
\newblock {\em Finance and Stochastics}, 15(1):57--8, 2011.

\bibitem{pen11c}
T.~Pennanen.
\newblock Convex duality in stochastic optimization and mathematical finance.
\newblock {\em Mathematics of Operations Research}, to appear.

\bibitem{pp10}
T.~Pennanen and I.~Penner.
\newblock Hedging of claims with physical delivery under convex transaction
  costs.
\newblock {\em SIAM Journal on Financial Mathematics}, 1:158--178, 2010.

\bibitem{rs5}
Mikl{\'o}s R{\'a}sonyi and Lukasz Stettner.
\newblock On utility maximization in discrete-time financial market models.
\newblock {\em Ann. Appl. Probab.}, 15(2):1367--1395, 2005.

\bibitem{roc70a}
R.~T. Rockafellar.
\newblock {\em Convex analysis}.
\newblock Princeton Mathematical Series, No. 28. Princeton University Press,
  Princeton, N.J., 1970.

\bibitem{roc74}
R.~T. Rockafellar.
\newblock {\em Conjugate duality and optimization}.
\newblock Society for Industrial and Applied Mathematics, Philadelphia, Pa.,
  1974.

\bibitem{rw76}
R.~T. Rockafellar and R.~J.-B. Wets.
\newblock Nonanticipativity and {$L^1$}-martingales in stochastic optimization
  problems.
\newblock {\em Math. Programming Stud.}, (6):170--187, 1976.
\newblock Stochastic systems: modeling, identification and optimization, II
  (Proc. Sympos., Univ Kentucky, Lexington, Ky., 1975).

\bibitem{rw98}
R.~T. Rockafellar and R.~J.-B. Wets.
\newblock {\em Variational analysis}, volume 317 of {\em Grundlehren der
  Mathematischen Wissenschaften [Fundamental Principles of Mathematical
  Sciences]}.
\newblock Springer-Verlag, Berlin, 1998.

\bibitem{shi96}
A.~N. Shiryaev.
\newblock {\em Probability}, volume~95 of {\em Graduate Texts in Mathematics}.
\newblock Springer-Verlag, New York, second edition, 1996.
\newblock Translated from the first (1980) Russian edition by R. P. Boas.

\bibitem{thi81}
Lionel Thibault.
\newblock Esp\'erances conditionnelles d'int\'egrandes semi-continus.
\newblock {\em Ann. Inst. H. Poincar\'e Sect. B (N.S.)}, 17(4):337--350, 1981.

\bibitem{tru91}
A.~Truffert.
\newblock Conditional expectation of integrands and random sets.
\newblock {\em Ann. Oper. Res.}, 30(1-4):117--156, 1991.
\newblock Stochastic programming, Part I (Ann Arbor, MI, 1989).

\bibitem{val75}
M.~Valadier.
\newblock Convex integrands on {S}ouslin locally convex spaces.
\newblock {\em Pacific J. Math.}, 59(1):267--276, 1975.

\end{thebibliography}

\documentclass{article}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{amsthm}

\def\reals{\mathbb{R}}

\def\uball{\mathbb{B}}
\def\ereals{\overline{\mathbb{R}}}

\def\comp{\raise 1pt \hbox{$\scriptstyle\circ$}}

\def\argmin{\mathop{\rm argmin}\limits}

\def\minimize{\mathop{\rm minimize}\limits}
\def\maximize{\mathop{\rm maximize}\limits}

\def\st{\mathop{\rm subject\ to}}

\def\dom{\mathop{\rm dom}\nolimits}

\def\lev{\mathop{\rm lev}\nolimits}

\def\upto{{\raise 1pt \hbox{$\scriptstyle \,\nearrow\,$}}}
\def\downto{{\raise 1pt \hbox{$\scriptstyle \,\searrow\,$}}}

\def\co{\mathop{\rm co}}

\def\FF{(\F_t)_{t=0}^T}

\def\A{{\cal A}}
\def\B{{\cal B}}
\def\C{{\cal C}}
\def\D{{\cal D}}

\def\F{{\cal F}}
\def\G{{\cal G}}

\def\L{{\cal L}}

\def\N{{\cal N}}

\def\U{{\cal U}}

\def\Y{{\cal Y}}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{example}{Example}
\newtheorem{remark}{Remark}
\theoremstyle{definition}

\title{Stochastic programs without duality gaps}
\author{Teemu Pennanen \and Ari-Pekka Perkki\"o}

\begin{document}

\maketitle

\begin{abstract}
This paper studies dynamic stochastic optimization problems parametrized by a random variable. Such problems arise in many applications in operations research and mathematical finance. We give sufficient conditions for the existence of solutions and the absence of a duality gap. Our proof uses extended dynamic programming equations, whose validity is established under new relaxed conditions that generalize certain no-arbitrage conditions from mathematical finance.
\end{abstract}

\section{Introduction}

Let $(\Omega,\F,P)$ be a probability space with a filtration $\FF$ (an increasing sequence of sub-sigma-algebras of $\F$) and consider the dynamic stochastic optimization problem
\begin{equation}\label{p}\tag{P}
\minimize\quad Ef(x(\omega),u(\omega),\omega)\quad\text{over $x\in\N$},
\end{equation}
where, for given integers $n_t$ and $m$
\[
\N = \{(x_t)_{t=0}^T\,|\,x_t\in L^0(\Omega,\F_t,P;\reals^{n_t})\},
\]
$u\in L^0(\Omega,\F,P;\reals^m)$ and $f$ is an extended real-valued convex {\em normal integrand} on $\reals^n\times\reals^m\times\Omega$, where $n=n_0+ \ldots +n_T$. Recall that $L^0(\Omega,\F_t,P;\reals^{n_t})$ denotes the space of equivalence classes of $\F_t$-measurable $\reals^{n_t}$-valued functions that coincide $P$-almost surely. That $f$ is a normal integrand, means that the set-valued mapping $\omega\mapsto\{(x,u,\alpha)\,|\,f(x,u,\omega)\le\alpha\}$ is closed-valued and $\F$-measurable; see e.g.\ \cite[Chapter~14]{rw98}. This implies that $f$ is $\F\otimes\B(\reals^n\times\reals^m)$-measurable (see \cite[Corollary~14.34]{rw98}), so that $\omega\mapsto f(x(\omega),u(\omega),\omega)$ is $\F$-measurable for every $x\in\N$. Throughout this paper, the expectation is defined for any measurable function by setting it equal to $+\infty$ unless the positive part is integrable. We will also assume that $\F$ as well as $\F_t$ for $t=0,\ldots,T$ are complete with respect to $P$\footnote{This allows us to use certain results on conditional expectations of integrands which are not necessarily normal in the general case. This is based on \cite[Corollary~14.34]{rw98}, which says that, when $\F$ is $P$-complete, then a function $f:\Omega\times\reals^n\times\reals^m\to\ereals$ is a normal integrand if and only if it is $\F\otimes\B(\reals^n\times\reals^m)$-measurable and $(x,u)\mapsto f(\omega,x,u)$ is lower semicontinuous for every $\omega$.}.

The measurable function $u$ may be interpreted as a parameter or a perturbation in a given stochastic optimization problem. It was shown in \cite{pen11c} that \eqref{p} covers many important problems in operations research and mathematical finance and how the conjugate duality framework of Rockafellar~\cite{roc74} allows for a unified treatment of many well-known duality frameworks. In that context, the lower-semicontinuity of the value function
\[
\varphi(u)=\inf_{x\in\N}Ef(x(\omega),u(\omega),\omega)
\]
over an appropriate space of measurable functions $u$ is equivalent to the absence of a duality gap; see \cite[Section~2]{pen11c} for a precise statement. In certain applications, most notably in mathematical finance, the objective in \eqref{p} lacks the inf-compactness properties required by the classical ``direct method'' of calculus of variations for establishing lower semicontinuity (and the existence of solutions). It was shown in \cite[Section~5]{pen11c} how certain measure theoretic techniques from mathematical finance can be combined with classical techniques of convex analysis to obtain the lower semicontinuity of $\varphi$. It is essential for this that the strategies $x\in\N$ are allowed to be general measurable functions not restricted to be e.g.\ integrable. The lower semicontinuity result given in \cite{pen11c}, however, applies to normal integrands $f$ that take only the values $0$ and $+\infty$. While that already covers some fundamental results in mathematical finance, as illustrated in \cite[Section~6]{pen11c}, it is far from satisfactory from the general point of view.

The main purpose of this paper is to establish the lower semicontinuity of $\varphi$ for more general normal integrands. This will be done in Section~\ref{sec:lsc}. Our proof extends that of \cite[Theorem~8]{pen11c}, which employs a recursive argument reminiscent of dynamic programming. We clarify this connection in Section~\ref{sec:dp} by generalizing the dynamic programming equations proposed by Rockafellar and Wets~\cite{rw76} for stochastic convex optimization. The dynamic programming equations were substantially generalized already by Evstigneev~\cite{evs76} who removed many of the assumptions made in \cite{rw76}, including convexity. We will show that in the convex case, the inf-compactness assumption made in both \cite{rw76} and \cite{evs76} can be replaced by weaker ``recession condition'' which subsumes, in particular, various no-arbitrage conditions used in mathematical finance. An early application of recession analysis to utility maximization in financial markets can be found in Bertsekas~\cite{ber74}. Section~\ref{sec:app} of this paper gives an application to an optimal consumption problem in illiquid markets.

\section{Dynamic programming}\label{sec:dp}

The purpose of this section is to extend the {\em dynamic programming} recursion of \cite[Section~3]{rw76} which generalizes the classical Bellman equation for convex stochastic optimization. We will use the notion of a conditional expectation of a normal integrand much as in \cite{evs76} where certain assumptions (convexity, nonanticipativity of the domain of $f$ and the boundedness of the strategies) of \cite{rw76} were relaxed. We show that, in the convex case, the inf-compactness assumption used in both \cite{rw76} and \cite{evs76} can be replaced by a milder condition on the directions of recession much like in the classical closedness results of finite-dimensional convex analysis; see \cite[Section~8]{roc70a}. In certain financial applications, the new condition turns out to be equivalent to the classical no-arbitrage condition.

Let $X$ be a nonnegative $\F$-measurable function and let $\G\subseteq\F$ be another sigma-algebra. Then, there is a $\G$-measurable nonnegative function $E^\G X$, unique up to sets of $P$-measure zero, such that
\begin{equation}\label{eq:ce}
E[\chi_AX]=E[\chi_A(E^\G X)]\quad\forall A\in\G,
\end{equation}
where $\chi_A$ denotes the characteristic function of $A$; see e.g.\ Shiryaev~\cite[II.7]{shi96}.
The function $E^\G X$ is called the $\G$-{\em conditional expectation} of $X$. For a general $\F$-measurable extended real-valued function $X$, we set
\[
E^\G X:=E^\G X_+-E^\G X_-,
\]
where again, the convention $\infty-\infty=\infty$ is used. It is easily checked that with the extended definition of the integral, \eqref{eq:ce} is then valid for any measurable function $X$. Our definition of conditional expectation extends \cite[Definition~II.7.1]{shi96}, which assumes that $\min\{E^\G X_+,E^\G X_-\}<\infty$ almost surely. Our choice of setting $\infty-\infty=\infty$ is not arbitrary but specifically directed towards minimization problems.

The $\G$-{\em conditional expectation} of a normal integrand $h$ is a $\G$-measurable normal integrand $E^\G h$ such that
\[
(E^\G h)(x(\omega),\omega)= E^\G [h(x(\cdot),\cdot)](\omega)\quad P\text{-a.s.}
\]
for all $x\in L^0(\Omega,\G,P;\reals^n)$. 
There are various conditions that guarantee the existence and uniqueness of a conditional expectation of a normal integrand; see e.g.\ Bismut~\cite{bis73}, Dynkin and Evstigneev~\cite{de76}, Castaing and Valadier~\cite[Section~VIII.9]{cv77}, Thibault~\cite{thi81}, Truffert~\cite{tru91} or Choirat, Hess and Seri~\cite{chs3}. The following suffices for the purposes of this paper.

\begin{lemma}\label{lem:ce}
Let $\G\subseteq\F$ be a sigma-algebra and assume that $h$ is an $\F$-normal integrand with an integrable lower bound i.e.\ an integrable function $m$ such that $h(x,\omega)\ge m(\omega)$ for every $x$ and $\omega$. Then $h$ has a well-defined conditional expectation $E^\G h$ which has the integrable lower bound $E^G m$.
\end{lemma}

\begin{proof}
The integrable lower bound implies, for example, the quasi-integrability condition of Thibault~\cite{thi81} as well as the condition of Choirat, Hess and Seri~\cite{chs3}, both of which give the existence and uniqueness of the conditional expectation. It follows from the monotonicity of the conditional expectation that if $f\ge m$ for an integrable function $m$, then $E^\G f\ge E^\G m$.
\end{proof}

We will study problem \eqref{p} for a fixed $u\in L^0(\omega,\F,P;\reals^m)$ so we will omit it from the notation and define 
\[
h(x,\omega)=f(x,u(\omega),\omega).
\]
By \cite[14.45(c)]{rw98}, $h$ is a normal integrand. The convexity of $f$ implies that of $h$. We will use the notation $E_t=E^{\F_t}$ and $x^t=(x_0,\ldots,x_t)$ and define extended real-valued functions $h_t,\tilde h_t:\reals^{n_1+\dots+n_t}\times\Omega\rightarrow\ereals$ recursively for $t=T,\ldots,0$ by
\begin{equation}\label{dp}
\begin{split}
\tilde h_T&=h,\\
h_t &= E_t\tilde h_t,\\
\tilde h_{t-1}(x^{t-1},\omega)&=\inf_{x_t\in\reals^{n_t}}h_t(x^{t-1},x_t,\omega).
\end{split}
\end{equation}
This is essentially the dynamic programming recursion introduced in \cite{rw76}. Our formulation with conditional expectations of normal integrands is closer to \cite{evs76}, where certain assumptions of \cite{rw76} were relaxed. In the above formulation, one does not separate the decision variables $x_t$ into  ``state'' and ``control'' like in the classical dynamic programming models; see e.g.\ \cite{ber76} and \cite{bs78}. A formulation closer to the classical dynamic programming equations will be given in Corollary~\ref{cor:bel} below. A recent application of dynamic programming to mathematical finance can be found in R{\'a}sonyi and Stettner~\cite[Section~5]{rs5}.

In order to ensure that $h_t$ and $\tilde h_t$ are well-defined it suffices to require that the function $h$ has an integrable lower bound and that $h(\cdot,\omega)$ is inf-compact (i.e.\ $\{x\in\reals^n\,|\,h(x,\omega)\le \alpha\}$ is compact for every $\alpha\in\reals$) for every $\omega\in\Omega$; see \cite[Theorem~5]{evs76}. In the convex case, the compactness assumption can be replaced by a weaker condition stated in terms of the {\em recession function} of $h$. If $\dom h(\cdot,\omega)$ is nonempty, then the recession function has the expression
\[
h^\infty(x,\omega)=\sup_{\lambda>0}\frac{h(\lambda x + \bar x,\omega)-h(\bar x,\omega)}{\lambda},
\]
which is independent of the choice of $\bar x\in\dom h(\cdot,\omega)$; see \cite[Theorem~8.5]{roc70a} or \cite[3.21]{rw98}. By \cite[Exercise~14.54(a)]{rw98}, the function $h^\infty$ is a convex normal integrand. If $h(\cdot,\omega)$ has an integrable lower bound, then $h^\infty(x,\omega)\ge 0$ for every $x\in\reals^n$ as is easily seen by letting $\lambda\to\infty$.

\begin{lemma}\label{lem:ip}
Assume that $h_t$ is a normal integrand and that the set-valued mapping
\[
N_t(\omega)=\{x_t\in\reals^{n_t}\,|\,h_t^\infty(x^t,\omega)\le 0,\ x^{t-1}=0\}
\]
is linear-valued. Then $\tilde h_{t-1}$ is a normal integrand with
\[
\tilde h^\infty_{t-1}(x^{t-1},\omega) = \inf_{x_t\in\reals^{n_t}}h^\infty_t(x^{t-1},x_t,\omega).
\]
Moreover, given an $x\in\N$, there is an $\F_t$-measurable $\bar x_t$ such that $\bar x_t(\omega)\perp N_t(\omega)$ and
\[
\tilde h_{t-1}(x^{t-1}(\omega),\omega) = h_t(x^{t-1}(\omega),\bar x_t
(\omega),\omega).
\]
\end{lemma}

\begin{proof}
By \cite[Theorem~9.2]{roc70a}, the linearity condition implies that the infimum in the definition of $\tilde h_{t-1}$ is attained and that $\tilde h_{t-1}(\cdot,\omega)$ is a lower semicontinuous convex function with
\[
\tilde h^\infty_{t-1}(x^{t-1},\omega)=\inf_{x_t\in\reals^{n_t}}h^\infty_t(x^{t-1},x_t,\omega).
\]
By \cite[Proposition~14.47]{rw98}, the lower semicontinuity implies that $\tilde h_{t-1}$ is an $\F_t$-measurable convex normal integrand. By \cite[Proposition~14.45(c)]{rw98}, the function $p(x,\omega):=h_t(x^{t-1}(\omega),x,\omega)$ is then also an $\F_t$-measurable normal integrand so, by \cite[Theorem~14.37]{rw98}, there is an $\F_t$-measurable $\bar x_t$ that attains the minimum for every $\omega$. By \cite[Corollary~8.6.1]{roc70a}, the value of $h_t(x^{t-1}(\omega),x,\omega)$ does not change if we replace $\bar x_t(\omega)$ by its projection to the orthogonal complement of $N_t(\omega)$. By \cite[Exercise~14.17]{rw98}, such a projection preserves measurability.
\end{proof}



It is clear that if $h_t$ has an integrable lower bound, then so will $\tilde h_{t-1}$. Applying Lemmas~\ref{lem:ce} and \ref{lem:ip} recursively backwards for $t=T,\ldots,0$, we then see that if $h$ has an integrable lower bound, the functions $\tilde h_t$ and $h_t$ are well-defined for every $t$ provided that $N_t$ is linear-valued at each step. 

We now get the following refinement of the optimality conditions in \cite[Theorem~1]{rw76} and \cite[Theorems~1 and 2]{evs76} in the convex case.

\begin{theorem}\label{thm:dp}
Assume that $h$ has an integrable lower bound and that $N_t$ is linear-valued for $t=T,\ldots,0$. The functions $h_t$ are then well-defined normal integrands and we have for every $x\in\N$ that
\begin{equation}\label{ie}
Eh_t(x_t(\omega),\omega)\ge \inf\eqref{p}\quad t=0,\ldots,T.
\end{equation}
Optimal solutions $x\in\N$ exist and they are characterized by the condition
\[
x_t(\omega)\in\argmin_{x_t}h_t(x^{t-1}(\omega),x_t,\omega)\quad P\text{-a.s.}\quad t=0,\ldots,T.
\]
which is equivalent to having equalities in \eqref{ie}. Moreover, there is an optimal solution $x\in\N$ such that $x_t\perp N_t$ for every $t=0,\ldots,T$.
\end{theorem}

\begin{proof}
As noted above, a recursive application of Lemmas~\ref{lem:ce} and~\ref{lem:ip} imply that the functions $h_t$ and $\tilde h_t$ are well-defined normal integrands. Given an $x\in\N$, the law of iterated expectations (see e.g.\ Shiryaev~\cite[Section~II.7]{shi96}) gives
\[
Eh_t(x^t(\omega),\omega) \ge E\tilde h_{t-1}(x^{t-1}(\omega),\omega) = Eh_{t-1}(x^{t-1}(\omega),\omega)\quad t=1,\ldots,T.
\]
Thus,
\[
Eh(x(\omega),\omega) = Eh_T(x^T(\omega),\omega) \ge Eh_0(x^0(\omega),\omega) \ge E\inf_{x_0\in\reals^{n_0}}h_0(x_0,\omega),
\]
where the inequalities hold as equalities if and only if
\[
h_t(x^t(\omega),\omega)=\tilde h_{t-1}(x^{t-1}(\omega),\omega)\quad P\text{-a.s.}\quad t=0,\ldots,T.
\]
The existence of such an $x\in\N$ with $x_t\perp N_t$ follows by applying Lemma~\ref{lem:ip} recursively for $t=0,\ldots,T$.
\end{proof}

When the normal integrand $h$ has a separable structure, the dynamic programming equations \eqref{dp} can be written in a more familiar form.

\begin{corollary}[Bellman equations]\label{cor:bel}
Assume that
\[
h(x,\omega) = \sum_{t=0}^T k_t(x_{t-1},x_t,\omega)
\]
for some fixed initial state $x_{-1}$ and $\F_t$-measurable normal integrands $h_t$ with integrable lower bounds. Consider the functions $V_t:\reals^{n_t}\times\Omega\to\ereals$ given by
\begin{equation}\label{bellman}
\begin{split}
V_T(x_T,\omega) &= 0,\\
\tilde V_{t-1}(x_{t-1},\omega) &= \inf_{x_t\in\reals^{n_t}}\{k_t(x_{t-1},x_t,\omega) + V_t(x_t,\omega)\},\\
V_{t-1} &= E_{t-1}\tilde V_{t-1}
\end{split}
\end{equation}
and assume that the set-valued mappings
\[
N_t(\omega) = \{x_t\in\reals^{n_t}\,|\, k_t^\infty(0,x_t,\omega) + V_t^\infty(x_t,\omega)\le 0\}
\]
are linear-valued for each $t=T,\ldots,0$. The functions $V_t$ are then well-defined normal integrands and we have for every $x\in\N$ that
\begin{equation}\label{ieb}
E\left[\sum_{s=0}^tk_s(x_{s-1}(\omega),x_s(\omega),\omega) + V_t(x_t(\omega),\omega)\right] \ge \inf\eqref{p}\quad t=0,\ldots,T.
\end{equation}
Optimal solutions $x\in\N$ exist and they are characterized by the condition
\[
x_t(\omega)\in\argmin_{x_t\in\reals^{n_t}}\{k_t(x_{t-1}(\omega),x_t,\omega) + V_t(x_t,\omega)\}\quad P\text{-a.s.}\quad t=0,\ldots,T,
\]
which is equivalent to having equalities in \eqref{ieb}. Moreover, there is an optimal solution $x\in\N$ such that $x_t\perp N_t$ for every $t=0,\ldots,T$.
\end{corollary}

\begin{proof}
By Theorem~\ref{thm:dp}, it suffices to show that
\begin{equation}\label{Qq}
h_t(x^t,\omega) = \sum_{s=0}^tk_s(x_{s-1},x_s,\omega) + V_t(x_t,\omega)
\end{equation}
for every $t=0,\ldots,T$. For $t=T$, \eqref{Qq} is obvious since $V_T=0$ by definition. Assuming that \eqref{Qq} holds for $t$, we get
\begin{align*}
\tilde h_{t-1}(x^{t-1},\omega) &= \inf_{x_t\in\reals^{n_t}} h_t(x^{t-1},x_t,\omega)\\
&= \sum_{s=0}^{t-1}k_s(x_{s-1},x_s,\omega) + \inf_{x_t\in\reals^{n_t}}\{k_t(x_{t-1},x_t,\omega) + V_t(x_t,\omega)\}\\
&= \sum_{s=0}^{t-1}k_s(x_{s-1},x_s,\omega) + \tilde V_{t-1}(x_{t-1},\omega)
\end{align*}
and then, since for $s=0,\ldots,t-1$, $k_s$ is $\F_{t-1}$-measurable,
\begin{align*}
h_{t-1}(x^{t-1},\omega) &= \sum_{s=0}^{t-1}k_s(x_{s-1},x_s,\omega) + V_{t-1}(x_{t-1},\omega),
\end{align*}
where, by Lemma~\ref{lem:ce}, $V_t$ is a well-defined normal integrand when $N_t$ is linear-valued.
\end{proof}

The rest of this section is devoted to the study of the linearity condition in Theorem~\ref{thm:dp}. Recall that a {\em $\G$-measurable selector} of an $\reals^n$-valued set-valued mapping $C$ is a $\G$-measurable function $x$ such that $x(\omega)\in C(\omega)$ almost surely.

\begin{lemma}\label{lem:rce}
Let $\G\subseteq\F$ be a sigma-algebra and assume that $h$ is an $\F$-normal integrand with an integrable lower bound. If there is an $\bar x\in L^0(\Omega,\G,P;\reals^n)$ such that $Eh(\bar x(\omega),\omega)$ is finite, then $(E^\G h)^\infty = E^\G h^\infty$ and the level sets 
\begin{align*}
\lev_0 h^\infty(\omega) &= \{x\in\reals^n\,|\, h^\infty(x,\omega)\le 0\},\\
\lev_0 (E^\G h)^\infty(\omega) &= \{x\in\reals^n\,|\, (E^\G h)^\infty(x,\omega)\le 0\}
\end{align*}
have the same $\G$-measurable selectors.
\end{lemma}

\begin{proof}
By \cite[Exercise~14.54]{rw98}, $h^\infty$ is a well-defined $\F$-normal integrand. Moreover, the lower bound on $h$ implies that $h^\infty$ is nonnegative. By Lemma~\ref{lem:ce}, $E^\G h$ and $E^\G h^\infty$ are thus well-defined. To show that the latter is the recession function of the former, let $x\in L^0(\Omega,\G,P;\reals^n)$ and $A\in\G$. Convexity of $h$ implies that the difference quotient
\[
\frac{h(\bar x(\omega) +\lambda x(\omega),\omega)-h(\bar x(\omega),\omega)}{\lambda}
\]
is increasing in $\lambda$ for every $\omega$; see e.g.\ \cite[Theorem~23.1]{roc70a}. The lower bound on $h$ and the integrability of $h(\bar x(\cdot),\cdot)$ thus imply that, for $\lambda\ge 1$, the quotients are minorized by a fixed integrable function. Monotone convergence theorem then gives for every $A\in\G$
\begin{align*}
E [1_A h^\infty(x)] &= E [1_A \lim_{\lambda\nearrow\infty} (h(\bar x+\lambda x)-h(\bar x))/\lambda]\\
&=\lim_{\lambda\nearrow\infty} E[1_A (h(\bar x+\lambda x)-h(\bar x))/\lambda]\\
&=\lim_{\lambda\nearrow\infty} E[1_A ((E^\G h)(\bar x+\lambda x)- (E^\G h)(\bar x))/\lambda]\\
&=E[1_A\lim_{\lambda\nearrow\infty} ((E^\G h)(\bar x+\lambda x)- (E^\G h)(\bar x))/\lambda]\\
&=E[1_A (E^\G h)^\infty(x)],
\end{align*}
which means that $(E^\G h)^\infty$ is the conditional expectation of $h^\infty$.

To prove the last claim, let $x\in L^0(\Omega,\G,P;\reals^n)$. By the first claim and the definition of a conditional integrand,
\[
(E^\G h^\infty)(x(\cdot),\cdot)=E^\G h^\infty(x(\cdot),\cdot).
\]
We have $h^\infty(x(\omega),\omega)\le 0$ almost surely if and only if $E^\G h^\infty(x(\cdot),\cdot)\le~0$ almost surely, since $h^\infty\geq 0$.
\end{proof}

\begin{remark}\label{rem:rec}
Consider the parametric problem \eqref{p} and assume that $u\in L^0(\Omega,\F,P;\reals^m)$ is such that $h(\cdot,\omega)=f(\cdot,u(\omega),\omega)$ is proper. We then have
\[
h^\infty(x,\omega)=f^\infty(x,0,\omega),
\]
where $f^\infty(\cdot,\cdot,\omega)$ is the recession function of $f(\cdot,\cdot,\omega)$. It follows that, as soon as they are well defined, the recession functions $\tilde h_t^\infty$ and $h_t^\infty$ and thus, the mappings $N_t$ are independent of the choice of $u\in\dom\varphi$. Indeed, $u\in\dom\varphi$ implies that there is an $x\in\N$ such that $h(x(\omega),\omega)=f(x(\omega),u(\omega),\omega)<\infty$ almost surely. Recursive application of Lemmas~\ref{lem:ip} and \ref{lem:rce} then shows that $h_t$ and $\tilde h_t$ can be expressed in terms of $h^\infty$, which is independent of $u\in\dom\varphi$.
\end{remark}

The following result shows that the linearity condition of Theorem~\ref{thm:dp} can be stated in terms of the original normal integrand $h$ directly. In the proof, we will denote the set of $\G$-measurable selectors of a set-valued mapping $C$ by $L^0(\G;C)$. We will also use the fact that if $C$ is closed-valued and $\G$-measurable, then it is almost surely linear-valued if and only if the set of its measurable selectors is a linear space. This follows easily by considering the Castaing representation of $C$; see e.g.~\cite[Theorem~14.5]{rw98}.

\begin{lemma}\label{lem:lin}
Assume that $h$ has an integrable lower bound and that $E h(\bar x(\omega),\omega)<\infty$ for some $\bar x \in\N$. Then $h_t$ is well-defined and $N_t$ is linear-valued for $t=T,\ldots,0$ if and only if 
\[
\L=\{x\in \N|\, h^\infty(x(\omega),\omega)\le 0 \text{ a.s.}\}
\]
is a linear space. If $x\in\L$ is such that $x^{t-1}=0$ then $x_t\in N_t$ almost surely.
\end{lemma}

\begin{proof}
Redefining $h(x,\omega):=h(x-\bar x(\omega),\omega)$, we may assume that $\bar x=0$. Indeed, such a translation amounts to translating the functions $\tilde h_t$ and $h_t$ accordingly and it does not affect the recession functions $\tilde h_t^\infty$ and $h_t^\infty$. We proceed by induction on $T$. When $T=0$, Lemma~\ref{lem:rce} gives
\begin{align*}
\L&= \{x\in\N|\, h_T^\infty(x(\omega),\omega)\le 0 \text{ a.s.}\} = L^0(\F_T;N_T).
\end{align*}
Since $N_T$ is $\F_T$-measurable, the linearity of $\L$ is equivalent to $N_T$ being linear-valued. Let now $T$ be arbitrary and assume that the claim holds for every $(T-1)$-period model.

If $\L$ is linear then $\L'=\{x\in \N|\, x_0=0,\ h^\infty(x(\omega),\omega)\le 0 \text{ a.s.}\}$ is linear as well. Applying the induction hypothesis to the $(T-1)$-period model obtained by fixing $x_0\equiv0$, we get that $N_t$ is linear for $t=T,\ldots,1$. Applying Lemmas~\ref{lem:ce} and \ref{lem:ip} backwards for $s=T,\ldots,1$, we then see that $h_0$ is well defined. Lemmas~\ref{lem:rce} and \ref{lem:ip} give
\begin{align}
L^0(\F_0;N_0) &= \{x_0\in L^0(\F_0)\,|\, h_0^\infty(x_0(\omega),\omega)\le 0\text{ a.s.}\}\nonumber\\
&= \{x_0\in L^0(\F_0)\,|\, \tilde h_0^\infty(x_0(\omega),\omega)\le 0\text{ a.s.}\}\nonumber\\
&= \{x_0\in L^0(\F_0)\,|\, \inf_{x_1}h_1^\infty(x_0(\omega),x_1,\omega)\le 0\text{ a.s.}\}\nonumber\\
&= \{x_0\in L^0(\F_0)\,|\, \exists\tilde x\in\N:\ \tilde x_0=x_0,\ h_1^\infty(\tilde x^1(\omega),\omega)\le 0\text{ a.s.}\}\nonumber,
\end{align}
where the last equality follows by applying the last part of Lemma~\ref{lem:ip} to the normal integrand $h^\infty$.
Repeating the argument for $t=1,\ldots,T$, we get
\begin{align}
L^0(\F_0;N_0) &= \{x_0\in L^0(\F_0)\,|\, \exists\tilde x\in\N: \tilde x_0=x_0,\ h_T^\infty(\tilde x(\omega),\omega)\le 0 \text{ a.s.}\}\nonumber\\
&= \{x_0\in L^0(\F_0)\,|\, \exists\tilde x\in\N: \tilde x_0=x_0,\ h^\infty(\tilde x(\omega),\omega)\le 0 \text{ a.s.}\}\nonumber\\
&= \{x_0\in L^0(\F_0)\,|\, \exists\tilde x\in\L: \tilde x_0=x_0\}.\label{n0}
\end{align}
The linearity of $\L$ thus implies that of $L^0(\F_0;N_0)$ which is equivalent to $N_0$ being linear-valued.

Assume now that $N_t$ is linear-valued for $t=T,\ldots,0$ and let $x\in\L$. Expression \eqref{n0} for $L^0(\F_0;N_0)$ is again valid so, by linearity of $N_0$, there is an $\tilde x\in\L$ with $\tilde x_0=-x_0$. Since $h^\infty$ is sublinear, $\L$ is a cone, so that $x+\tilde x\in\L$. Since $x_0+\tilde x_0=0$, we also have $x+\tilde x\in\L'$. Since, by the induction assumption, $\L'$ is linear and since $\L'\subseteq\L$, we get $-x-\tilde x\in\L$. Since $\L$ is a cone, we get $-x=\tilde x-x-\tilde x\in\L$. Thus, $\L$ is linear.

For $t=0$, the last claim follows directly from expression \eqref{n0}. The general case follows by applying this to the $(T-t)$-period model obtained by fixing $x^{t-1}\equiv 0$.
\end{proof}

When $h$ is the indicator function of a convex set, the linearity condition in Lemma~\ref{lem:lin} becomes the linearity condition of \cite[Theorem~8]{pen11c} which generalizes various no-arbitrage conditions that have been used in mathematical finance. The following example illustrates the situation in the classical perfectly liquid market model; see \cite{pen11c} for more general models.

\begin{example}[Superhedging in liquid markets]\label{ex:sh}
Let $S=(S_t)_{t=0}^T$ be an $\reals^d$-valued $\FF$-adapted stochastic process, $n_t=d$, $m=1$ and
\[
f(x,u,\omega) = 
\begin{cases}
0 & \text{if $\sum_{t=0}^{T-1}x_t\cdot\Delta S_t(\omega)\ge u$,}\\
+\infty & \text{otherwise}.
\end{cases}
\]
We get
\[
\varphi(u)=\inf_{x\in\N} I_f(x,u) =
\begin{cases}
0 & \text{if $×u\in\C$},\\
+\infty & \text{otherwise},
\end{cases}
\]
where $\C = \{u\in L^0\,|\,\exists x\in\N:\ \sum_{t=0}^{T-1}x_t\cdot\Delta S_t\ge u\}$. In the classical perfectly liquid model of financial markets, where $S$ gives the unit prices of the ``risky assets'' and $x_t$ is the portfolio held over $(t,t+1]$, the set $\C$ consists of the {\em contingent claims} that can be {\em superhedged} without a cost; see e.g.~\cite[Section~6.4]{ds6}. Since $f$ is a closed positively homogeneous function, we have $f^\infty=f$ and
\[
\{x\in \N\,|\, f^\infty(x(\omega),0,\omega)\le 0 \text{ a.s.}\}=\{x\in \N|\, \sum_{t=0}^{T-1}x_t\cdot\Delta S_t\ge 0\}.
\]
This set is linear, and thus, the function $h(x,\omega)=f(x,u(\omega),\omega)$ satisfies the linearity condition in Lemma~\ref{lem:lin}, if and only if the price process $S$ satisfies the {\em no-arbitrage} condition. 
\end{example}

The following simple example goes beyond indicator functions and also of inf-compact integrands considered in \cite{rw76,evs76}.

\begin{example}[Variance optimal hedging]\label{ex:voh}
Let $S=(S_t)_{t=0}^T$ be an $\reals^d$-valued $\FF$-adapted stochastic process, $u\in L^0(\Omega,\F,P;\reals)$ and consider the problem of minimizing
\[
E(V_0 + \sum_{t=0}^{T-1} z_t\cdot\Delta S_{t+1} - u)^2
\]
over $V_0\in\reals$ and $\F_t$-measurable $\reals^d$-valued functions $z_t$. This corresponds to \eqref{p} with $x_0=(z_0,V_0)$, $x_t=z_t$ for $t=1,\ldots,T$ and 
\[
f(x,u,\omega) = (V_0 + \sum_{t=0}^{T-1} z_t\cdot\Delta S_{t+1}(\omega) - u)^2.
\]
The above problem has been studied e.g.\ in F\"ollmer and Schied~\cite[Section~10.3]{fs4}, where $V_0$ is interpreted as an initial value of a self-financing trading strategy where $z_t$ is the portfolio of risky assets held over period $[t,t+1]$. By \cite[Theorem~9.4]{roc70a},
\[
f^\infty(x,u,\omega) = 
\begin{cases}
0 & \text{if $V_0 + \sum_{t=0}^{T-1} z_t\cdot\Delta S_{t+1}(\omega) - u=0$},\\
+\infty & \text{otherwise}.
\end{cases}
\]
By Remark~\ref{rem:rec}, the function $h(x,\omega)=f(x,u(\omega),\omega)$ then satisfies the linearity condition of Lemma~\ref{lem:lin}, so the optimal solution is attained. This should be compared with the existence results in \cite[Section~10.3]{fs4}, where it was assumed that $d=1$.
\end{example}

\section{Lower semicontinuity of the value function}\label{sec:lsc}

We now return to the parametrized problem \eqref{p}. Being the inf-projection of the convex integral functional
\[
I_f(x,u)=Ef(x(\omega),u(\omega),\omega),
\]
the value function 
\[
\varphi(u)=\inf_{x\in\N}Ef(x(\omega),u(\omega),\omega)
\]
is convex on $L^0(\Omega,F,P;\reals^m)$; see e.g.\ \cite[Theorem~1]{roc74}. Our aim is to give conditions under which $\varphi$ is lower semicontinuous on certain locally convex topological vector subspaces of $L^0(\Omega,F,P;\reals^m)$. The lower semicontinuity is equivalent to the absence of a duality gap in the duality framework of \cite{pen11c} (which is essentially an instance of the conjugate duality framework of Rockafellar~\cite{roc74}) which we now briefly recall (and slightly generalize).

Assume that $\U$ and $\Y$ are vector subspaces of $L^0(\Omega,F,P;\reals^m)$ in {\em separating duality} under the bilinear form
\[
\langle u,y\rangle = E[u(\omega)\cdot y(\omega)],
\]
i.e.\ that $E[u(\omega)\cdot y(\omega)]$ is finite for every $u\in\U$ and $y\in\Y$ and that for every nonzero $u\in\U$ (resp.\ $y\in\Y$), there is at least one $y\in\Y$ (resp.\ $u\in\U$) such that $\langle u,y\rangle\ne 0$. The special case $\U=L^p$ and $\Y=L^q$ was studied in \cite{pen11c}. The weakest and the strongest locally convex topologies on $\U$ compatible with the pairing will be denoted by $\sigma(\U,\Y)$ and $\tau(\U,\Y)$, respectively. Since the value function $\varphi$ is convex, we have by the classical separation argument, that $\varphi$ is lower semicontinuous with respect to $\sigma(\U,\Y)$ if it is merely lower semicontinuous with respect to $\tau(\U,\Y)$. When $\U=L^p$ and $\Y=L^q$ for $p\in[1,\infty)$, $\tau(\U,\Y)$ is simply the norm topology on $\U$ and $\sigma(\U,\Y)$ the weak topology. A general treatment of topological spaces in separating duality can be found e.g.\ in Kelley and Namioka~\cite{kn76}.

The {\em Lagrangian} associated with \eqref{p} is the extended real-valued function
\[
L(x,y) = \inf_{u\in\U}\{I_f(x,u)-\langle u,y\rangle\}
\]
on $\N\times\Y$. The Lagrangian is convex in $x$ and concave in $y$. The {\em dual objective} is the extended real-valued function on $\Y$ defined by 
\[
g(y) = \inf_{x\in\N}L(x,y).
\]
The basic duality result \cite[Theorem~7]{roc74} says, in particular, that $g = -\varphi^*$. When $\varphi$ is lower semicontinuous and proper, the biconjugate theorem (see e.g.\ \cite[Theorem~5]{roc74}) then gives the dual representation
\begin{equation}\label{bc}
\varphi(u) = \sup\{\langle u,y\rangle + g(y)\}.
\end{equation}
It was shown in \cite{pen11c} that this abstract result is behind many duality frameworks in stochastic optimization and mathematical finance.

It was assumed in \cite{pen11c} that $\U=L^p$ and $\Y=L^q$, but the main result \cite[Theorem~3]{pen11c} remains valid as long as the space $\U$ is {\em decomposable} in the sense that
\[
\chi_Au+\chi_{\Omega\setminus A}u'\in\U
\]
whenever $A\in\F$, $u\in\U$ and $u'\in L^\infty(\Omega,\F,P;\reals^m)$. Indeed, the decomposability property allows the use of the interchange rule for minimization and integration (see \cite[Theorem~14.60]{rw98}), which suffices for the proof of \cite[Theorem~3]{pen11c}. We also note that decomposability of the spaces $\U$ and $\Y$ implies that the separation property holds automatically for the bilinear form defined above; see \cite[Lemma~6]{val75}. 
Moreover, we have the following relations for relative topologies.

\begin{lemma}\label{lem:rel}
If $\U$ and $\Y$ are decomposable, then $L^\infty\subseteq\U\subseteq L^1$ and
\begin{align*}
\sigma(L^1,L^\infty)|_\U&\subseteq\sigma(\U,\Y),\quad\sigma(\U,\Y)|_{L^\infty}\subseteq\sigma(L^\infty,L^1),\\
\tau(L^1,L^\infty)|_\U&\subseteq\tau(\U,\Y),\quad\tau(\U,\Y)|_{L^\infty}\subseteq\tau(L^\infty,L^1).
\end{align*}
\end{lemma}

\begin{proof}
By \cite[Lemme 1, p.??]{bis73}, $L^\infty\subseteq\U\subset L^1$ and $L^\infty\subseteq\Y\subseteq L^1$ which give the relations for the $\sigma$-topologies. Since, by symmetry, analogous relations are valid for the $\sigma$-topologies on $\Y$, we have that $\sigma(L^\infty,L^1)$-compact subsets of $L^\infty$ are $\sigma(\Y,\U)$-compact. Since, by the Mackey-Arens theorem, $\tau(\U,\Y)$ is generated by the support functions of $\sigma(\Y,\U)$-compact sets, we get $\tau(L^1,L^\infty)|_\U\subseteq\tau(\U,\Y)$. The remaining inclusion is verified similarly.
\end{proof}

The traditional ``direct method'' for proving the lower semicontinuity would be to assume that the integral functional $I_f$ is uniformly inf-compact in $x$ with respect to an appropriate topology on $\N$. If the topology is strong enough to imply the almost sure convergence of a subsequence, the sequential lower semicontinuity can often be derived from Fatou's lemma. In certain applications, this purely topological argument fails because $I_f$ lacks an appropriate inf-compactness property in $x$. In convex problems, the following measure theoretic result can sometimes be used as a substitute for compactness.

\begin{lemma}[Koml\'os' theorem]\label{lem:kom}
Let $(x^\nu)_{\nu=1}^\infty$ be a sequence in $L^0(\Omega,\F,P;\reals^n)$ which is either
\begin{enumerate}
\item
bounded in $L^1$,
\item
almost surely bounded in the sense that 
\[
\sup_\nu|x^\nu(\omega)|<\infty\quad P\text{-a.s.}
\]
\end{enumerate}
Then there is a sequence of convex combinations $\bar x^\nu\in\co\{x^\mu\,|\,\mu\ge\nu\}$ that converges almost surely to an $\reals^n$-valued function.
\end{lemma}

\begin{proof}
See e.g.\ \cite{ds6} or \cite{ks9}.
\end{proof}

The following is our main result.

\begin{theorem}\label{thm:lsc}
Assume that there is a $y\in\Y$ and an $m\in L^1(\Omega,\F,P)$ such that for $P$-almost every $\omega$,
\[
f(x,u,\omega)\ge u\cdot y(\omega) +m(\omega)\quad \forall (x,u)\in\reals^n\times\reals^m 
\]
and that $\{x\in \N|\, f^\infty(x(\omega),0,\omega)\le 0 \text{ a.s.}\}$ is a linear space. Then
\[
\varphi(u) = \inf_{x\in\N}I_f(x,u)
\]
is lower semicontinuous on $\U$ and the infimum is attained for every $u\in\U$.
\end{theorem}

\begin{proof}
Let $h_u(x,\omega)=f(x,u(\omega),\omega)$. The lower bound on $f$ implies that $h_u$ has an integrable lower bound. As noted in Remark~\ref{rem:rec}, $h_u^\infty(x,\omega)=f^\infty(x,0,\omega)$ for every $u$, so the linearity condition on $f$ implies that $h_u$ satisfies the linearity conditions in Lemma~\ref{lem:lin}. By Theorem~\ref{thm:dp}, the infimum in $\varphi(u)= \inf_{x\in\N}I_f(x,u)$ is thus attained for every $u\in\U$ by an $x\in\N$ with $x_t(\omega)\perp N_t(\omega)$ almost surely.

For lower semicontinuity, it suffices to show that $\varphi$ is lower semicontinuous on the linear space $L^{1,y}=\{u\in L^1|\, |E[u\cdot y]|<\infty\}$ with respect to the norm $|| u||_{L^{1,y}}=E|u|+ |E[u\cdot y]|$. Indeed, since $y\in\Y$, we have $\U\subseteq L^{1,y}$ and, by Lemma~\ref{lem:rel}, the norm $||\cdot||_{L^{1,y}}$ is continuous on $\tau(\U,\Y)$, which means that $\tau(\U,\Y)$ is stronger than the norm topology restricted to $\U$. Since $L^{1,y}$ is a normed space, it suffices to prove sequential lower semicontinuity, which means that for any $\gamma\in\reals$ and for any sequence $(u^\nu)_{\nu=1}^\infty$ such that 
\[
\varphi(u^\nu)\le\gamma
\]
and $u^\nu\to u$ in $L^{1,y}$, we have $\varphi(u)\le\gamma$. We will prove this by establishing the existence of an $x\in\N$ such that $I_f(x,u)\le\gamma$.

As observed at the beginning of the proof, there is for every $\nu$ an $x^\nu\in\N$ such that $x^\nu_t\perp N_t$ and
\[
I_f(x^\nu,u^\nu)\le \gamma.
\]
Moreover, the mappings $N_t$ are independent of $u^\nu$; see Remark~\ref{rem:rec}.
Since $u^\nu$ converges in $L^{1,y}$, the lower bound on $f$ implies that the negative parts of the functions $\omega\mapsto f(x^\nu(\omega),u^\nu(\omega),\omega)$ are bounded in $L^1$. Since $I_f(x^\nu,u^\nu)\le \gamma$, the positive parts must be bounded as well. Thus, by Lemma~\ref{lem:kom}, there is a sequence of convex combinations 
\[
\phi^\nu(\omega):=\sum_{\mu=\nu}^\infty\alpha^{\nu,\mu}f(x^\mu(\omega),u^\mu(\omega),\omega)
\]
that converges almost surely to a real-valued measurable function. In particular, the function $\phi(\omega):=\sup_\nu\phi^\nu(\omega)$ is almost surely finite. Defining 
\[
(\bar x^\nu,\bar u^\nu) = \sum_{\mu=\nu}^\infty \alpha^{\nu,\mu}(x^\mu,u^\mu)
\]
we have by convexity that 
\[
f(\bar x^\nu(\omega),\bar u^\nu(\omega),\omega)\le\phi^\nu(\omega)\le\phi(\omega)\quad P\text{-a.s.}
\]
and $I_f(\bar x^\nu,\bar u^\nu)\le \gamma$. Moreover, we still have $\bar x^\nu_t\in N_t^\perp$ almost surely and $\bar u^\nu\to u$ in the $L^{1,y}$-norm. 

Passing to a subsequence if necessary, we may assume that $\bar u^\nu\to u$ almost surely, so that the measurable function $\rho(\omega) := \sup_\nu|\bar u^\nu(\omega)|$ is almost surely finite. Each $(\bar x^\nu,\bar u^\nu)$ then belongs to the set
\[
\C=\{(x,u)\in\N\times L^0\,|\, (x,u)\in C\ \text{a.s.}\},
\]
where $C(\omega) = \{(x,u)\,|\, x_t\in N_t^\perp(\omega),\ u\in\rho(\omega)\uball,\ f(x,u,\omega)\le\phi(\omega)\}$. We will now apply \cite[Theorem~6]{pen11c}, which says that the sequence $(\bar x^\nu,\bar u^\nu)_{\nu=1}^\infty$ is almost surely bounded if  
\begin{equation}\label{zero}
\{(x,u)\in\N\times L^0\,|\, (x,u)\in C^\infty\ \text{a.s.}\}=\{(0,0)\}.
\end{equation}
By Corollary~8.3.3 and Theorem~8.7 of \cite{roc70a}, 
\[
C^\infty(\omega)=\{(x,0)\,|\, x_t\in N_t^\perp(\omega),\ f^\infty(x,0,\omega)\le 0\}.
\]
If $x\in\N$ is such that $f^\infty(x(\omega),0,\omega)\le 0$ then, by the last part of Lemma~\ref{lem:lin}, we have $x_0\in N_0$. The condition $x_0\in N_0^\perp$ then implies that $x_0=0$. Repeating the argument for $t=1,\ldots,T$ gives \eqref{zero} so $(\bar x^\nu,\bar u^\nu)_{\nu=1}^\infty$ is almost surely bounded.

By Lemma~\ref{lem:kom}, there is a sequence $(\hat x^\nu,\hat u^\nu)_{\nu=1}^\infty$ of convex combinations of $(\bar x^\nu,\bar u^\nu)_{\nu=1}^\infty$ that converges almost surely to a point $(x,\hat u)$, where necessarily $\hat u=u$ since $\bar u^\nu\to u$ almost surely. We still have $\hat u^\nu\to u$ in the $L^{1,y}$-norm and, by convexity, $I_f(\hat x^\nu,\hat u^\nu)\le\gamma$. By Fatou's lemma,
\begin{multline*}
E[f(x(\omega),u(\omega),\omega)-y(\omega)\cdot u(\omega)-m(\omega)]\\
\le\liminf_{\nu\to\infty} E[f(\hat x^\nu(\omega),\hat u^\nu(\omega),\omega)-y(\omega)\cdot\hat u^\nu(\omega)-m(\omega)],
\end{multline*}
where $E[y(\omega)\cdot\hat u^\nu(\omega)]\to E[y(\omega)\cdot u(\omega)]$, by the $L^{1,y}$-convergence, so that
\[
I_f(x,
u)\le \liminf_{\nu\to\infty}I_f(\hat x^\nu,\hat u^\nu)\le \gamma,
\]
which completes the proof.
\end{proof}

\section{An application to mathematical finance}\label{sec:app}

We will illustrate Theorem~\ref{thm:lsc} on the optimal consumption problem considered in \cite[Section~5]{pen11c}. The problem is set in a generalization of the market model of Kabanov~\cite{kab99}, where a finite number $d$ of securities is traded over finite discrete time $t=0,\ldots,T$. At each time $t$ and state $\omega\in\Omega$, the market is described by two closed convex sets, $C_t(\omega)$ and $D_t(\omega)$, both of which contain the origin. The set $C_t(\omega)$ consists of the portfolios that are freely available in the market and $D_t(\omega)$ consists of the portfolios that the investor is allowed to hold over the period $[t, t+1)$. For each $t$, the sets $C_t$ and $D_t$ are assumed to be $\F_t$-measurable. 

Consider the problem
\begin{align}\label{ocp}
\begin{split}
\maximize_{(z,c)\in\N}\quad & E\sum_{t=0}^T U_t(c_t)\\
\st \quad  & z_t-z_{t-1} + c_t \in C_t,\ z_t\in D_t\quad P\text{-a.s.}\ t=0,\ldots,T,
\end{split}
\end{align}
where $z_{-1}:=0$, $D_T(\omega):=\{0\}$ and $-U_t$ is a convex $\F_t$-measurable normal integrand on $\reals^d\times\Omega$. This models an optimal consumption problem where at each time $t$ and stage $\omega$ we can consume some of the assets and update the existing portfolio $z_{t-1}$. The combined process $(z,c)$ is required to be {\em self-financing} in the sense that the sum of the portfolio update $\Delta z_t:=z_t-z_{t-1}$ and the consumption vector $c_t$ has to be freely available in the market, i.e.\ it belongs to $C_t(\omega)$. In addition, the portfolio constraint $z_t(\omega)\in D_t(\omega)$ is required to hold almost surely at each time. Problem \eqref{ocp} generalizes the classical optimal consumption problem where the numeraire asset is consumed in a perfectly liquid market model (see Examples~\ref{ex:sh} and \ref{ex:voh}). A general treatment of the continuous-time model can be found in Karatzas and \u{Z}itkovi\'c~\cite{kz3}.

Defining
\[
\C = \{c\in\A\,|\, \Delta z_t + c_t \in C_t,\ z_t\in D_t\quad P\text{-a.s.}\ t=0,\ldots,T\},
\]
where $\A$ denotes the set of $\reals^d$-valued adapted processes (so that $\N=\A\times\A$), we can write problem \eqref{ocp} compactly as
\begin{equation}\label{ocpc}
\maximize_{c\in\A}\quad E\sum_{t=0}^T U_t(c_t)\quad\text{over $c\in\C$}.
\end{equation}
The set $\C$ can be interpreted as the set of consumption processes that can be super-replicated without a cost in the market given by the pair $(C,D)$; compare with the definition of the set $\C$ in Example~\ref{ex:sh}.

In order to dualize the problem, we embed it in the general duality framework with, $x_t=(z_t,c_t)$, $u=(u_t)_{t=0}^T$ and 
\begin{align*}
f(x,u,\omega)= 
\begin{cases}
-\sum_{t=0}^T U_t(c_t,\omega) & \text{if }\Delta z_t + c_t +u_t \in C_t(\omega),\, z_t\in D_t(\omega)\\
+\infty &\text{otherwise}.
\end{cases}
\end{align*}
Here $u_t\in\reals^d$ so that the dimension of $u$ equals $m=(T+1)d$. The Lagrangian integrand becomes
\begin{align*}
&\ l(x,y,\omega) = \inf_{u\in\reals^m}\{f(x,u,\omega)-u\cdot y\}\\
&=
\inf_{u\in\reals^m}\{-\sum_{t=0}^T[U_t(c_t,\omega)+u_t\cdot y_t]\, | \, \Delta z_t + c_t +u_t \in C_t(\omega),\ z_t\in D_t(\omega)\}\\
&=
\inf_{\tilde u\in\reals^m}\{-\sum_{t=0}^T [U_t(c_t,\omega)+(\tilde u_t-\Delta z_t-c_t)\cdot y_t]\, | \, \tilde u_t \in C_t(\omega),\ z_t\in D_t(\omega)\}\\
&= 
\begin{cases}
-\sum_{t=0}^T [U_t(c_t,\omega)+\sigma_{C_t(\omega)}(y_t) - (\Delta z_t+c_t)\cdot y_t] & \text{if $z_t\in D_t(\omega)$}\\
+\infty & \text{otherwise}
\end{cases}\\
&= 
\begin{cases}
-\sum_{t=0}^T [U_t(c_t,\omega) + \sigma_{C_t(\omega)}(y_t) + z_t\cdot\Delta y_{t+1} - c_t\cdot y_t] & \text{if $z_t\in D_t(\omega)$}\\
+\infty & \text{otherwise},
\end{cases}
\end{align*}
where $\sigma_{C_t(\omega)}$ denotes the support function of $C_t(\omega)$. In the last equality we have used the ``integration by parts'' formula
\[
\sum_{t=0}^T\Delta z_t\cdot y_t = -\sum_{t=0}^Tz_t\cdot\Delta y_{t+1}
\]
where $y_{T+1}:=0$. 

Recall that, by Lemma~\ref{lem:rel}, $\Y\subset L^1$. On the other hand, by \cite[Theorem 3]{pen11c}\footnote{Theorem 3 of \cite{pen11c} is stated for the case $\U=L^p$ and $\Y=L^q$, but its proof goes through in the general case without a change.},
\begin{align*}
g(y)=\inf_{x\in\N^\infty} El(x(\omega),y(\omega),\omega),
\end{align*}
where $\N^\infty=\N\cap L^\infty$. We can then use the law of iterated expectations (see e.g.\ \cite[Section~II.7]{shi96}) and the interchange rule for integration and minimization (see e.g.\ \cite[Theorem~14.60]{rw98}) to write the dual objective as
\begin{align*}
g(y) &= \inf_{(c,z)\in\N^\infty}\left\{ E\sum_{t=0}^T[-U_t(c_t)-\sigma_{C_t}(y_t) \right.\\
&\left. \qquad\qquad\qquad\qquad {} - z_t\cdot E_{t}\Delta y_{t+1}+c_t\cdot E_t y_t] \left.\right|\,  z_t\in D_t\text{ a.s.} \vphantom{E\sum_{t=0}^T}\right\}\\
&= E\sum_{t=0}^T \inf_{c_t,z_t\in\reals^d}\left\{-U_t(c_t,\omega)-\sigma_{C_t(\omega)}(y_t(\omega)) \right.\\
 &\left. \qquad\qquad\qquad\quad {}-z_t\cdot(E_t\Delta y_{t+1})(\omega)+c_t\cdot(E_ty_t)(\omega)|\, z_t\in D_t(\omega)\right\}\\
&= E\sum_{t=0}^T [U^*_t(E_ty_t) -\sigma_{C_t}(y_t)-\sigma_{D_t}(E_t\Delta y_{t+1})].
\end{align*}
where 
\[
U_t^*(y,\omega) = \inf_{c\in\reals^d}\{c\cdot y-U_t(c,\omega)\}
\]
is the conjugate of $U_t$ in the concave sense.

When $C_t(\omega)$ and $D_t(\omega)$ are convex cones, we have
\[
g(y)=
\begin{cases}
E\sum_{t=0}^T U^*_t(E_ty_t) & \text{if $y\in\D$},\\
-\infty & \text{otherwise},
\end{cases}
\]
where
\[
\D = \{y\in\Y\,|\,y_t\in C_t^*,\ E_t\Delta y_t\in D_t^*\},
\]
where $C_t^*(\omega)$ and $D_t^*(\omega)$ are the polar cones of $C_t(\omega)$ and $D_t(\omega)$, respectively. The dual problem can then be written as
\begin{equation}\label{ocpd}
\maximize\quad E\sum_{t=0}^TU^*_t(y_t)\quad\text{{\rm over}}\ y\in\D
\end{equation}
in symmetry with the primal problem \eqref{ocpc}. The $\FF$-adapted elements of $\D$ are called {\em consistent price systems} for the market model $(C,D)$; see \cite[Example~4.2]{pen11c}. It follows from Jensen's inequality that the value of the dual objective does not decrease when replacing a general $y\in\Y$ by its $\FF$-adapted projection; see \cite[Example~3.6]{pen11c}.

In \cite{pen11c}, the lower semicontinuity of the value function associated with the optimal consumption problem was left open so it could not be claimed that the optimal values of \eqref{ocpc} and \eqref{ocpd} are equal. With the help of Theorem~\ref{thm:lsc}, we can now derive simple sufficient conditions. We will assume that the utility function $U_t$ satisfies the growth condition
\begin{equation}\label{gc}
U_t^\infty(c,\omega)=\begin{cases}
0 & \text{if $c\in\reals_+^d$},\\
-\infty & \text{otherwise},
\end{cases}
\end{equation}
for every $\omega$. When $d=1$ and $U_t(\cdot,\omega)$ is smooth, this is equivalent to the conditions $\lim_{c\to\infty}U_t'(c,\omega)=0$ and $\lim_{c\to-\infty}U_t'(c,\omega)=+\infty$ which generalize the Inada conditions; see e.g.\ \cite{ks99}. We will say that the market model $(C,D)$ satisfies the condition of {\em no-scalable arbitrage} if 
\begin{equation}\label{nsa}
\{c\in\A_+\,|\,\exists z\in\A:\ \Delta z_t + c_t\in C^\infty_t,\ z_t\in D^\infty_t\}=\{0\},
\end{equation}
where $\A_+$ denotes the set of componentwise nonnegative claim processes. This condition is related to arbitrage opportunities that can be scaled up by arbitrarily large positive numbers; see \cite{pen11,pp10}.

\begin{theorem}\label{thm:ocp}
Assume that the optimal value of \eqref{ocp} or, equivalently, \eqref{ocpc} is less than $\infty$, that $U_t$ satisfy the growth condition \eqref{gc} and that there is an integrable function $m$ such that $U_t(c,\omega)\le m(\omega)$ almost surely for every $c\in\reals^d$ and $t=0,\ldots,T$. If the market model $(C,D)$ satisfies the no-scalable arbitrage condition \eqref{nsa} and if the set
\begin{align*}
&\{z\in\A\,|\, \Delta z_t \in C^\infty_t,\ z_t\in D^\infty_t\},
\end{align*}
is linear, then the optimal value of
\[
\maximize\quad E\sum_{t=0}^T U_t(c_t)\quad\text{over}\quad c\in\C-u
\]
is lower semicontinuous as a function of $u\in\U$ and the infimum is always attained. In particular, if $C$ and $D$ are conical, then the optimal value of the primal \eqref{ocpc} is the negative of the optimal value of the dual \eqref{ocpd}.
\end{theorem}

\begin{proof}
By \cite[Theorems~9.3 and 9.5]{roc70a},
\begin{align*}
f^\infty(x,u,\omega)=
\begin{cases}
-\sum_{t=0}^T U^\infty_t(c_t,\omega) & \text{if }\Delta z_t + c_t \in C^\infty_t(\omega),\, z_t\in D^\infty_t(\omega)\\
+\infty & \text{otherwise},
\end{cases}
\end{align*}
so, by Lemma~\ref{lem:lin}, the linearity condition of Theorem~\ref{thm:lsc} means that the set
\[
\L=\{(z,c)\in\N \, | \, \sum_{t=0}^TU_t^\infty(c_t)\ge 0,\ \Delta z_t + c_t\in C^\infty_t,\, z_t\in D^\infty_t\}
\]
is linear. Under the growth condition \eqref{gc}, 
\[
\L=\{(z,c)\in\N \, | \, c_t\ge 0,\ \Delta z_t + c_t\in C^\infty_t,\, z_t\in D^\infty_t\}
\]
If $(z,c)\in\L$, condition \eqref{nsa} implies that $c=0$, and then, by linearity of the set $\{z\in\A\,|\, \Delta z_t \in C^\infty_t,\ z_t\in D^\infty_t\}$, we have $(-z,-c)\in\L$. Since $\L$ is also a cone, it has to be a linear space.
\end{proof}

\begin{remark}\label{rem:rnsa}
The conclusions of Theorem~\ref{thm:ocp} remain valid if, instead of the growth condition \eqref{gc} and the no-scalable arbitrage condition \eqref{nsa}, we assume that the set
\begin{equation}\label{lin3}
\{c\in\A\,|\,\exists z\in\A:\ \sum_{t=0}^TU_t^\infty(c_t)\ge 0,\ \Delta z_t + c_t\in C^\infty_t,\ z_t\in D^\infty_t\},
\end{equation}
is linear. Indeed, if $\L$ is as in the above proof and $(z,c)\in\L$, condition \eqref{lin3} gives the existence of a $z^-\in\A$ such that $(z^-,-c)\in\L$. Since $\L$ is a cone, we get $(z+z^-,0)\in\L$ and then, the linearity condition of Theorem~\ref{thm:ocp} gives $(-(z+z^-),0)\in\L$. Since $-(z,c)=(-(z+z^-),0)+(z^-,-c)$, we get that $-(z,c)\in\L$, i.e., $\L$ is linear.
\end{remark}

\bibliographystyle{plain}
\bibliography{sp}

\end{document}